# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ 3 Teacher-Student Problems
[…]
@@
Here $f_{\mathcal{V}}$ is the “teacher,” and is assumed to be fixed, while $f_{\mathcal{W}}$ the “student.” The problem (9) is equivalent to low-rank tensor approximation, with respect to a possibly nonstandard norm.
@@ 3.1 Functional norms
We begin by describing functional norms of interest. To fix notation, consider $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ and $\mathcal{V}=(\beta,V)\in\mathbb{R}^{s}\times\mathbb{R}^{s\times n}$, and let $\tau({\mathcal{W}})=\sum_{i=1}^{r}\alpha_{i}w_{i}^{\otimes d}$ and $\tau({\mathcal{V}})=\sum_{i=1}^{s}\beta_{i}v_{i}^{\otimes d}$ be tensors in $\textup{Sym}^{d}(\mathbb{R}^{n})$ such that $f_{\mathcal{W}}(x)=\langle\tau(\mathcal{W}),x^{\otimes r}\rangle$ and $f_{\mathcal{V}}(x)=\langle\tau(\mathcal{V}),x^{\otimes r}\rangle$. We will focus our attention on the following two types of norms:
@@
2) Distribution-induced norms. Let $\mathcal{D}$ be a probability distribution over $\mathbb{R}^{d}$ (decaying fast enough). Then $\|f_{\mathcal{W}}-f_{\mathcal{V}}\|_{\mathcal{D}}^{2}\,:=\langle f_{\mathcal{W}}-f_{\mathcal{V}},f_{\mathcal{W}}-f_{\mathcal{V}}\rangle_{\mathcal{D}}$, where
@@
A special class of distribution-induced norms arise from discrete distributions $\mathcal{D}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$, where $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{n}$. This is the setting of empirical risk minimization, where $\|f_{\mathcal{W}}-f_{\mathcal{V}}\|_{ERM}^{2}:=\langle f_{\mathcal{W}}-f_{\mathcal{V}},f_{\mathcal{W}}-f_{\mathcal{V}}\rangle$ for
$\langle g,h\rangle_{ERM}\,:=\frac{1}{N}\sum_{i=1}^{N}g(x_{i})h(x_{i})\quad\textup{for all }g,h:\mathbb{R}^{d}\rightarrow\mathbb{R}.$ (12)
@@
The ERM inner product (12) defines a non-degenerate inner product over ${\rm Sym}^{r}(\mathbb{R}^{n})$ if and only if there exist no non-zero $n$-variate polynomial of degree $d$ that vanishes on $x_{1},\ldots,x_{N}$. This is generically true for $N\geq\binom{n+d-1}{d}$.
@@
Distribution-induced products can always be expressed in terms of the $2d$-th moments of $\mathcal{D}$. For $g,h$ homogeneous degree $d$ forms, written as $g(x)=\langle S,x^{\otimes d}\rangle_{F}$ and $h(x)=\langle T,x^{\otimes d}\rangle_{F}$ where $S,T\in\textup{Sym}^{d}(\mathbb{R}^{n})$, we have that
$\langle g,h\rangle_{\mathcal{D}}$ $=\mathbb{E}_{x\sim\mathcal{D}}[\langle S,x^{\otimes d}\rangle_{F}\langle T,x^{\otimes d}\rangle_{F}]$ (13)
$=\mathbb{E}_{x\sim\mathcal{D}}[\langle S\otimes T,x^{\otimes 2d}\rangle_{F}]$
$=\left\langle S\otimes T,\mathbb{E}_{x\sim\mathcal{D}}\,x^{\otimes 2d}\right\rangle_{F}$
$=\langle S\otimes T,{M}_{\mathcal{D},2d}\rangle_{F}$
$=\textup{vec}(S)^{\top}\textup{mat}({M}_{\mathcal{D},2d})\textup{vec}(T).$
[…]
@@
where $[n]:=\{1,\ldots,n\}$ and $S=(s_{i_{1},\ldots,i_{d}})$, $T=(t_{i_{1},\ldots,i_{d}})$ and $x=(x_{i})$. From this it follows that $\textup{mat}({M}_{\mathcal{D},2d})$ is never a diagonal matrix, and so $\langle g,h\rangle_{\mathcal{D}}\neq{\rm vec}(S)^{T}{\rm vec}(T)$ for all $\mathcal{D}$. ∎
@@
$\mathcal{D}$ with i.i.d. coordinates. Here $x\in\mathbb{R}^{d}$, with i.i.d. coordinates according to some probability distribution $\pi$ on $\mathbb{R}$. Equivalently, $\mathcal{D}=\pi^{\otimes d}$. Now for each $s\geq 0$, denote by $\mu_{\pi,s}$ the $s$-th moment of $\pi$, i.e., $\mu_{\pi,s}=\mathbb{E}_{y\sim\pi}[y^{s}]\in\mathbb{R}$. Also, given $I\in[n]^{2d}$, we let #$\#(I,s)$ be the number of elements of $[n]$ that occur exactly $s$ times in $I$. Then,
@@
Moments of distributions with independent but not necessarily identical coordinates are derived in [43, 3].
@@
The RHS is proportional to $\|A^{T}z\|_{2}^{2d}=\langle AA^{T},zz^{T}\rangle_{F}^{d}=\langle\textup{Sym}(\Sigma^{\otimes d}),z^{\otimes 2d}\rangle_{F}$, as desired. Formulas for the moment tensors of non-centered Gaussians may be found in [38].
Mixtures $\mathcal{D}$. Let $\mathcal{D}=\sum_{i=1}^{t}\gamma_{i}\mathcal{D}_{i}$ be a convex combination of distributions $D_{i}$. Then,
@@
Many properties of the moments of various mixture models are developed in [38, 43, 3].
@@
Assume $n=d=2$. A distribution-induced inner product on the space $2\times 2$ matrices is described by a $4\times 4$ matrix of $4$-th moments:
@@ 3.2 Critical points of the distance function
Let us now take a step back and describe some general aspects of the distance function of points to smooth manifolds that will be used to study the landscape of student-teacher problems for shallow polynomial networks. Due to the generality of the setup, results in this subsection may also be of independent interest.
@@
${\rm Crit}(L_{t,\Sigma}|_{X}):=\{s\in\text{Smooth}(X):dL_{t,\Sigma}(T_{s}X)=0\}.$ (20)
Since almost all functions are Morse functions, ${\rm Crit}(L_{t,\Sigma}|_{X})$ is typically a finite set. Using local parameterizations, we can also associate a Hessian signature to each critical point $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$. More precisely, fix a chart $p:U\rightarrow V$, where $U\subseteq\mathbb{R}^{d}$ and $V\subseteq X$ are open neighborhoods respectively of 00 and $s$, such that $p(0)=s$; then the signature and rank of $H(L\circ p)(0)$ where $H$ denotes the (usual) Hessian matrix does not depend on the choice of chart. We sometimes use $H_{L,X}$ to denote any Hessian matrix associated with a chart, focusing solely on its rank or signature.
@@
$Z_{X}:=\{(s,t,\Sigma)\in X\times\mathbb{R}^{n}\times\operatorname{PD}_{n}\colon s\in{\rm Crit}(L_{t,\Sigma}|_{X})\text{ s.t. }\det H_{L_{t,\Sigma}}(s)=0\},$
$F_{X}:=\{(t,\Sigma)\in\mathbb{R}^{n}\times\operatorname{PD}_{n}\colon\exists s\in X\text{ s.t. }(s,t,\Sigma)\in Z_{X}\},$
$F_{X,\Sigma}:=\{t\in\mathbb{R}^{n}\colon(t,\Sigma)\in F_{X}\}.$
The set $Z_{X}$ consists of triples $(s,t,\Sigma)$ of student, teacher, and inner product, such that $s$ is a degenerate critical point for $L_{t,\Sigma}|_{X}$. The other sets are derived from $Z_{X}$. Classically, $F_{X,\Sigma}$ is known as the focal locus of $X$ (typically with $\Sigma={\rm Id}_{n}$). When $X$ is an algebraic variety, its Zariski closure is also called the ED discriminant [16]. A point $t\in\mathbb{R}^{n}$ is called a focal point of $(X,s)$ of multiplicity $m\geq 1$ for the distance $\Sigma$ if $(s,t,\Sigma)\in Z_{X}$ and the Hessian $H_{L_{t,\Sigma}}(s)$ has nullity $m$. The following result is adapted from [37, Chapter 6].
@@
Consider a fixed $\Sigma\in\operatorname{PD}_{n}$ and let $N_{X,\Sigma}:=\{(s,v)\in X\times\mathbb{R}^{n}\colon\langle v,w\rangle_{\Sigma}=0\,\,\forall w\in T_{s}X\}$ be the total space of the normal bundle to $X$. A point $t\in\mathbb{R}^{n}$ is a focal point of $(X,s)$ of multiplicity $m$ if and only if the differential of the “endpoint map” $e_{X,\Sigma}:N_{X,\Sigma}\rightarrow\mathbb{R}^{n},(s,v)\mapsto s+v$ has nullity $m$ at $(s,t-s)$. In particular, $F_{X,\Sigma}={\rm Br}(e_{X,\Sigma})$, where Br denotes the branch locus (Section 2.1).
@@
Without loss of generality we assume $\Sigma={\rm Id}_{n}$. Let $x(u)\in X$ be a local parameterization of $X$ with $u\in\mathbb{R}^{d}$ and consider $n-d$ vector fields $w_{1}(u),\ldots,w_{n-d}(u)$ in $\mathbb{R}^{n}$ that form an orthonormal frame of the normal bundle to $X$. We parameterize $N_{X,\Sigma}$ as $(s,v)=\left(x(u),\sum_{m=1}^{n-d}\gamma_{m}w_{m}(u)\right)$ with coordinates $(u,\gamma)\in\mathbb{R}^{d}\times\mathbb{R}^{n-d}$. If $t=s+v$, we have that
[…]
@@
$\frac{\partial^{2}L}{\partial u_{i}\partial u_{j}}=2\frac{\partial}{\partial u_{i}}\left(\frac{\partial x}{\partial u_{j}}\cdot(x-t)\right)=2\left(\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{i}}-\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot(t-x)\right).$ (24)
Thus, if $t=x(u)+\sum_{m=1}^{n-d}\gamma_{m}w_{m}(u)$, then (23) and (24) are the same up to the factor of $2$, and the nullity of the differential of the endpoint map $e_{X,\Sigma}$ at $(s,t-s)$ is the same as that of the Hessian $H_{L_{t,\Sigma}}(s)$. ∎
@@
As shown in the proof of Proposition 23, if $t=s+\gamma v$ where $v$ is a unit vector orthogonal to $X$, then the Hessian of the distance function $H_{L_{t,\Sigma}}(s)$ can be written as $G-\gamma(v\cdot E)$ with $G_{ij}:=\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{i}}$ and $(v\cdot E)_{ij}:=v\cdot\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}$. The bilinear forms $G$ and $(v\cdot E)$ are called respectively the first fundamental form and the second fundamental form in the direction of $v$. If local coordinates are chosen so that $G={\rm Id}_{d}$, then the eigenvalues of $(v\cdot E)$ are called the principal curvatures of $X$ at $s$ in the normal direction of $v$. In this setting, it is easy to see that $t$ is a focal point if and only if $\gamma=1/k$ where $k$ is a principal curvature.
@@
Let $X\subset\mathbb{R}^{n}$ be a smooth manifold, and assume that $X$ is either compact or is a cone in $\mathbb{R}^{n}$ (without the vertex). If $(t,\Sigma)\in(\mathbb{R}^{n}\times\operatorname{PD}_{n})\setminus F_{X}$, then there exists an open neighborhood $\mathcal{U}\subseteq\mathbb{R}^{n}\times\operatorname{PD}_{n}$ of $(t,\Sigma)$ such that:
@@
i) Consider an arbitrary critical point $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$. According to Proposition 23, the fact that $(t,\Sigma)\not\in F_{X}$ implies that the endpoint map $e_{X,\Sigma}:(s,v)\mapsto s+v$ is a local diffeomorphim at $(s,t-s)$. Moreover, the same is true for the “extended” endpoint map $e_{X}:(s,v,\Sigma)\mapsto(s+v,\Sigma)$, since the last component is the identity. It now follows from the Inverse Function Theorem that $e_{X}$ can be locally inverted in a neighborhood of $(s,t-s,\Sigma)$, so $s$ can be smoothly parameterized by $t$ and $\Sigma$. Finally, we observe that if $X$ is compact or a cone, then there exists a neighborhood $\mathcal{U}$ of $(t,\Sigma)$ such that the number of critical points ${\rm Crit}(L_{t^{\prime},\Sigma^{\prime}}|_{X})$ is locally constant for $(t^{\prime},\Sigma^{\prime})\in\mathcal{U}$, and all critical points for $(t^{\prime},\Sigma^{\prime})\in\mathcal{U}$ can be smoothly parameterized by $(t^{\prime},\Sigma^{\prime})$. To see this, we note that under our assumptions we can find a neighborhood $\mathcal{V}$ of $(t,\Sigma)$ such that $e_{X}^{-1}(\mathcal{V})\subset\mathcal{K}$ where $\mathcal{K}$ is compact (in other words, critical students are bounded): this is clear when $X$ is compact, while if $X$ is a cone it is true since $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$ implies $\langle t-s,s\rangle_{\Sigma}=0$ so $\|s\|_{\Sigma}^{2}\leq\|s\|_{\Sigma}^{2}+\|t-s\|^{2}_{\Sigma}=\|t\|^{2}_{\Sigma}$. Now since $(t,\Sigma)$ is a regular value of the extended endpoint map, $e_{X}^{-1}(t,\Sigma)=\{(s_{1},v_{1},\Sigma),\ldots,(s_{k},v_{k},\Sigma)\}$ is a finite set (it is discrete and contained in $\mathcal{K}$). If we consider pairwise disjoint neighborhoods $U_{1},\ldots,U_{k}\subset N_{X}$ of $(s_{i},v_{i},\Sigma)$ that are mapped diffeomorphically to $V_{1},\ldots,V_{k}\subset\mathbb{R}^{n}\times\operatorname{PD}_{n}$, we can then let $\mathcal{U}=V_{1}\cap\ldots\cap V_{k}\setminus e_{X}(\mathcal{K}\setminus(U_{1}\cup\ldots\cup U_{k}))$.
@@
ii) Since by construction, $\mathcal{U}\subset(\mathbb{R}^{n}\times\operatorname{PD}_{n})\setminus F_{X}$ we have $\det H_{L_{t^{\prime},\Sigma}}(s_{i})\neq 0$ for all critical students $s_{i}$, so the result follows the from continuity of the eigenvalues of the Hessian. ∎

### v2

@@ 3 Teacher-Student Problems
Section overview. We study teacher–student training problems, where a student network approximates a fixed teacher function via a loss over the function space. For shallow polynomial networks, this setting corresponds to low-rank tensor approximation under various norms, including distribution-induced norms determined by the training data. Such norms capture how the data distribution shapes optimization. We examine their effect on the optimization landscape through moment tensors and analyze critical points using a geometric framework for distance functions to manifolds. This leads to the notion of data discriminants, which reveal how changes in the data can induce qualitative transitions in the landscape.
We consider loss functions of the form
@@
where $f_{\mathcal{V}}$ is the fixed “teacher” and $f_{\mathcal{W}}$ is the trainable “student.” Minimizing $L(\mathcal{W})$ corresponds to approximating one shallow network by another, and when $s$ is smaller than $r$ is equivalent to a low-rank tensor approximation problem under a potentially nonstandard norm that may encode the training input distribution.
@@ 3.1 Functional norms
We begin by describing functional norms of interest. To fix notation, consider $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ and $\mathcal{V}=(\beta,V)\in\mathbb{R}^{s}\times\mathbb{R}^{s\times n}$, and let $\tau({\mathcal{W}})=\sum_{i=1}^{r}\alpha_{i}w_{i}^{\otimes d}$ and $\tau({\mathcal{V}})=\sum_{i=1}^{s}\beta_{i}v_{i}^{\otimes d}$ be tensors in $\textup{Sym}^{d}(\mathbb{R}^{n})$ such that $f_{\mathcal{W}}(x)=\langle\tau(\mathcal{W}),x^{\otimes r}\rangle_{{F}}$ and $f_{\mathcal{V}}(x)=\langle\tau(\mathcal{V}),x^{\otimes r}\rangle_{{F}}$. We will focus our attention on the following two types of norms:
@@
2) Distribution-induced norms. Let $\mathcal{D}$ be a probability distribution over $\mathbb{R}^{{n}}$ (decaying fast enough). Then $\|f_{\mathcal{W}}-f_{\mathcal{V}}\|_{\mathcal{D}}^{2}\,:=\langle f_{\mathcal{W}}-f_{\mathcal{V}},f_{\mathcal{W}}-f_{\mathcal{V}}\rangle_{\mathcal{D}}$, where
@@
A special class of distribution-induced norms arise from discrete distributions $\mathcal{D}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$, where $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{n}$. This is the setting of empirical risk minimization, where $\|f_{\mathcal{W}}-f_{\mathcal{V}}\|_{ERM}^{2}:=\langle f_{\mathcal{W}}-f_{\mathcal{V}},f_{\mathcal{W}}-f_{\mathcal{V}}\rangle_{{ERM}}$ for
$\langle g,h\rangle_{ERM}\,:=\frac{1}{N}\sum_{i=1}^{N}g(x_{i})h(x_{i})\quad\textup{for all }g,h:\mathbb{R}^{{n}}\rightarrow\mathbb{R}.$ (12)
@@
The ERM inner product (12) defines a non-degenerate inner product over ${\rm Sym}^{{d}}(\mathbb{R}^{n})$ if and only if there exists no non-zero $n$-variate polynomial of degree $d$ that vanishes on $x_{1},\ldots,x_{N}$. This is generically true for $N\geq\binom{n+d-1}{d}$.
@@
Distribution-induced inner products can always be expressed in terms of the $2d$-th moments of $\mathcal{D}$.
Proposition 21.
Let $g,h$ be homogeneous degree-$d$ forms, written as $g(x)=\langle S,x^{\otimes d}\rangle_{F}$ and $h(x)=\langle T,x^{\otimes d}\rangle_{F}$ for $S,T\in\mathrm{Sym}^{d}(\mathbb{R}^{n})$. The distribution-induced product (11) can be expressed in terms of the $2d$-th moment tensor of $\mathcal{D}$ as
$\langle g,h\rangle_{\mathcal{D}}=\langle S\otimes T,M_{\mathcal{D},2d}\rangle_{F}=\mathrm{vec}(S)^{\top}\,\mathrm{mat}(M_{\mathcal{D},2d})\,\mathrm{vec}(T),$
where $M_{\mathcal{D},2d}:=\mathbb{E}_{x\sim\mathcal{D}}[x^{\otimes 2d}]\in\mathrm{Sym}^{2d}(\mathbb{R}^{n})$ is the $2d$-th moment tensor, and $\mathrm{vec}$ and $\mathrm{mat}$ denote the vectorization and matricization operators with respect to lexicographic ordering of indices (mapping $\textup{Sym}^{d}(\mathbb{R}^{n})$ to $\mathbb{R}^{n^{d}}$ and $\textup{Sym}^{2d}(\mathbb{R}^{n})$ to $\mathbb{R}^{n^{d}\times n^{d}}$, respectively).
Proof.
Starting from the definition, we have
$\langle g,h\rangle_{\mathcal{D}}=\mathbb{E}_{x\sim\mathcal{D}}\big[\langle S,x^{\otimes d}\rangle_{F}\langle T,x^{\otimes d}\rangle_{F}\big]=\left\langle S\otimes T,\mathbb{E}_{x\sim\mathcal{D}}\,x^{\otimes 2d}\right\rangle_{F},$ (13)
which is precisely $\langle S\otimes T,M_{\mathcal{D},2d}\rangle_{F}$. The matricized form follows by applying $\mathrm{vec}$ to $S$ and $T$. ∎
Note that the representation of distribution-induced inner product in terms of moments makes no distinction between discrete and continuous distributions; in the latter case, we simply consider discrete moments.
@@
where $n]:=\{1,\ldots,n\}$ and $S=(s_{i_{1},\ldots,i_{d}})$, $T=(t_{i_{1},\ldots,i_{d}})$ and $x=(x_{i})$. From this it follows that $\textup{mat}({M}_{\mathcal{D},2d})$ is never a diagonal matrix (for example, consider the entries where $(i_{1},\ldots,i_{d})=(j_{1},\ldots,j_{d})$). Thus, $\langle g,h\rangle_{\mathcal{D}}\neq{\rm vec}(S)^{T}{\rm vec}(T)$ for all $\mathcal{D}$. ∎
@@
$\mathcal{D}$ with i.i.d. coordinates. Here $x\in\mathbb{R}^{{n}}$, with i.i.d. coordinates according to some probability distribution $\pi$ on $\mathbb{R}$. Equivalently, $\mathcal{D}=\pi^{\otimes{n}}$. Now for each $s\geq 0$, denote by $\mu_{\pi,s}$ the $s$-th moment of $\pi$, i.e., $\mu_{\pi,s}=\mathbb{E}_{y\sim\pi}[y^{s}]\in\mathbb{R}$. Also, given $I\in[n]^{2d}$, we let #(I,s)\#(I,s) be the number of elements of $[n]$ that occur exactly $s$ times in $I$. Then,
@@
Moments of distributions with independent but not necessarily identical coordinates are derived in [zhang2023moment, alexandr2023moment].
@@
The RHS is proportional to $\|A^{T}z\|_{2}^{2d}=\langle AA^{T},zz^{T}\rangle_{F}^{d}=\langle\textup{{sym}}(\Sigma^{\otimes d}),z^{\otimes 2d}\rangle_{F}$, as desired. Formulas for the moment tensors of non-centered Gaussians may be found in [pereira2022tensor].
Mixtures $\mathcal{D}$. Let $\mathcal{D}=\sum_{i=1}^{t}\gamma_{i}\mathcal{D}_{i}$ be a convex combination of distributions $\mathcal{D}_{i}$. Then,
@@
Many properties of the moments of various mixture models are developed in [pereira2022tensor, zhang2023moment, alexandr2023moment].
@@
Assume $n=d=2$. A distribution-induced inner product on the space of $2\times 2$ matrices is described by a $4\times 4$ matrix of $4$-th moments:
@@ 3.2 Critical points of the distance function
Let us now take a step back and describe some general aspects of the distance function of points to smooth manifolds that will be used to study the landscape of teacher-student problems for shallow polynomial networks. Due to the generality of the setup, results in this subsection may also be of independent interest.
@@
${\rm Crit}(L_{t,\Sigma}|_{X}):=\{s\in{X_{\rm reg}}:dL_{t,\Sigma}(T_{s}X)=0\},$ (20)
where $X_{\rm reg}$ denotes the set of smooth points on $X$, and $dL_{t,\Sigma}(T_{s}X)=0$ means that the differential of $L_{t,\Sigma}$ at $s$ vanishes in all directions tangent to $X$. Since almost all functions are Morse functions, ${\rm Crit}(L_{t,\Sigma}|_{X})$ is typically a finite set. Using local parameterizations, we can also associate a Hessian signature to each critical point $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$. More precisely, fix a chart $p:U\rightarrow V$, where $U\subseteq\mathbb{R}^{d}$ and $V\subseteq X$ are open neighborhoods respectively of 0 and $s$, such that $p(0)=s$; then the signature and rank of $H(L\circ p)(0)$ where $H$ denotes the (usual) Hessian matrix does not depend on the choice of chart. We sometimes use $H_{L,X}$ to denote any Hessian matrix associated with a chart, focusing solely on its rank or signature.
@@
$Z_{X}=\{(s,t,\Sigma)\in X\times\mathbb{R}^{n}\times\operatorname{PD}_{n}\colon s\in{\rm Crit}(L_{t,\Sigma}|_{X})\text{ s.t. }\det H_{L_{t,\Sigma}}(s)=0\},$
$F_{X}=\{(t,\Sigma)\in\mathbb{R}^{n}\times\operatorname{PD}_{n}\colon\exists s\in X\text{ s.t. }(s,t,\Sigma)\in Z_{X}\},$
$F_{X,\Sigma}=\{t\in\mathbb{R}^{n}\colon(t,\Sigma)\in F_{X}\}.$
The set $Z_{X}$ consists of triples $(s,t,\Sigma)$ of student, teacher, and inner product, such that $s$ is a degenerate critical point for $L_{t,\Sigma}|_{X}$. The other sets are derived from $Z_{X}$. Classically, $F_{X,\Sigma}$ is known as the focal locus of $X$ (typically with $\Sigma$ the identity). When $X$ is an algebraic variety, its Zariski closure is also called the ED discriminant [draisma2013]. A point $t\in\mathbb{R}^{n}$ is called a focal point of $(X,s)$ of multiplicity $m\geq 1$ for the distance $\Sigma$ if $(s,t,\Sigma)\in Z_{X}$ and the Hessian $H_{L_{t,\Sigma}}(s)$ has nullity $m$. The following result is adapted from [milnor2016morse, Chapter 6].
@@
Consider a fixed $\Sigma\in\operatorname{PD}_{n}$ and let $N_{X,\Sigma}:=\{(s,v)\in X\times\mathbb{R}^{n}\colon\langle v,w\rangle_{\Sigma}=0\,\,\forall w\in T_{s}X\}$ be the total space of the normal bundle to $X$. A point $t\in\mathbb{R}^{n}$ is a focal point of $(X,s)$ of multiplicity $m$ if and only if the differential of the “endpoint map” $e_{X,\Sigma}:N_{X,\Sigma}\rightarrow\mathbb{R}^{n},(s,v)\mapsto s+v$ has nullity $m$ at (s,ts)(s,t-s). In particular, $F_{X,\Sigma}={\rm Br}(e_{X,\Sigma})$, where Br denotes the branch locus (Section 2.1).
@@
Without loss of generality we assume $\Sigma={\rm Id}_{n}$. Let $x(u)\in X$ be a local parameterization of $X$ with $u\in\mathbb{R}^{d}$ and consider ndn-d vector fields $w_{1}(u),\ldots,w_{n-d}(u)$ in $\mathbb{R}^{n}$ that form an orthonormal frame of the normal bundle to $X$. We parameterize $N_{X,\Sigma}$ as $(s,v)=\left(x(u),\sum_{m=1}^{n-d}\gamma_{m}w_{m}(u)\right)$ with coordinates $(u,\gamma)\in\mathbb{R}^{d}\times\mathbb{R}^{n-d}$. If $t=s+v$, we have that
$\left\{\begin{aligned} &\frac{\partial t}{\partial u_{i}}=\frac{\partial x}{\partial u_{i}}+\sum_{m=1}^{n-d}\gamma_{{m}}\frac{\partial w_{m}}{\partial u_{{i}}},\qquad&&i=1,\ldots d,\\[5.69046pt] &\frac{\partial t}{\partial\gamma_{l}}=w_{l},&&l=1,\ldots,n-d.\end{aligned}\right$. (21)
@@
$\frac{\partial^{2}L}{\partial u_{i}\partial u_{j}}=2\frac{\partial}{\partial u_{i}}\left(\frac{\partial x}{\partial u_{j}}\cdot(x-t)\right)=2\left(\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{{j}}}-\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot(t-x)\right).$ (24)
Thus, if $t=x(u)+\sum_{m=1}^{n-d}\gamma_{m}w_{m}(u)$, then (23) and (24) are the same up to the factor of $2$, and the nullity of the differential of the endpoint map $e_{X,\Sigma}$ at (s,ts)(s,t-s) is the same as that of the Hessian $H_{L_{t,\Sigma}}(s)$. ∎
@@
As shown in the proof of Proposition 24, if $t=s+\gamma v$ where $v$ is a unit vector orthogonal to $X$, then the Hessian of the distance function $H_{L_{t,\Sigma}}(s)$ can be written as $G-\gamma(v\cdot E)$ with $G_{ij}:=\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{{j}}}$ and $(v\cdot E)_{ij}:=v\cdot\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}$. The bilinear forms $G$ and $(v\cdot E)$ are called respectively the first fundamental form and the second fundamental form in the direction of $v$. If local coordinates are chosen so that $G={\rm I}_{d}$, then the eigenvalues of $(v\cdot E)$ are called the principal curvatures of $X$ at $s$ in the normal direction of $v$. In this setting, it is easy to see that $t$ is a focal point if and only if $\gamma=1/k$ where $k$ is a principal curvature.
@@
Let $X\subset\mathbb{R}^{n}$ be a smooth manifold, and assume that $X$ is either compact or is a cone in $\mathbb{R}^{n}$ without the vertex. If $(t,\Sigma)\in(\mathbb{R}^{n}\times\operatorname{PD}_{n})\setminus F_{X}$, then there exists an open neighborhood $\mathcal{U}\subseteq\mathbb{R}^{n}\times\operatorname{PD}_{n}$ of $(t,\Sigma)$ such that:
@@
i) Consider an arbitrary critical point $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$. According to Proposition 24, the fact that $(t,\Sigma)\not\in F_{X}$ implies that the endpoint map $e_{X,\Sigma}:(s,v)\mapsto s+v$ is a local diffeomorphism at (s,ts)(s,t-s). Moreover, the same is true for the “extended” endpoint map $e_{X}:(s,v,\Sigma)\mapsto(s+v,\Sigma)$, since the last component is the identity. It now follows from the Inverse Function Theorem that $e_{X}$ can be locally inverted in a neighborhood of $(s,t-s,\Sigma)$, so $s$ can be smoothly parameterized by $t$ and $\Sigma$. Finally, we observe that if $X$ is compact or a cone, then there exists a neighborhood $\mathcal{U}$ of $(t,\Sigma)$ such that the number of critical points ${\rm Crit}(L_{t^{\prime},\Sigma^{\prime}}|_{X})$ is locally constant for $(t^{\prime},\Sigma^{\prime})\in\mathcal{U}$, and all critical points for $(t^{\prime},\Sigma^{\prime})\in\mathcal{U}$ can be smoothly parameterized by $(t^{\prime},\Sigma^{\prime})$. To see this, we note that under our assumptions we can find a neighborhood $\mathcal{V}$ of $(t,\Sigma)$ such that $e_{X}^{-1}(\mathcal{V})\subset\mathcal{K}$ where $\mathcal{K}$ is compact (in other words, critical students are bounded): this is clear when $X$ is compact, while if $X$ is a cone it is true since $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$ implies $\langle t-s,s\rangle_{\Sigma}=0$ so $\|s\|_{\Sigma}^{2}\leq\|s\|_{\Sigma}^{2}+\|t-s\|^{2}_{\Sigma}=\|t\|^{2}_{\Sigma}$. Now since $(t,\Sigma)$ is a regular value of the extended endpoint map, $e_{X}^{-1}(t,\Sigma)=\{(s_{1},v_{1},\Sigma),\ldots,(s_{k},v_{k},\Sigma)\}$ is a finite set (it is discrete and contained in $\mathcal{K}$). If we consider pairwise disjoint neighborhoods $U_{1},\ldots,U_{k}\subset N_{X}$ of $(s_{i},v_{i},\Sigma)$ that are mapped diffeomorphically to $V_{1},\ldots,V_{k}\subset\mathbb{R}^{n}\times\operatorname{PD}_{n}$, we can then let $\mathcal{U}=V_{1}\cap\ldots\cap V_{k}\setminus e_{X}(\mathcal{K}\setminus(U_{1}\cup\ldots\cup U_{k}))$.
@@
ii) Since by construction, $\mathcal{U}\subset(\mathbb{R}^{n}\times\operatorname{PD}_{n})\setminus F_{X}$ we have $\det H_{L_{t^{\prime},\Sigma}}(s_{i})\neq 0$ for all critical students $s_{i}$, so the result follows the from the continuity of the eigenvalues of the Hessian. ∎